OpenStudy (anonymous):

Which description matches the solution to this equation: x^4- 81 = 0 A. all four solutions are real B. two solutions are real; two are imaginary C. all four solutions are imaginary D. three solutions are real; one solution is imaginary

OpenStudy (anonymous):

Two are real, two are imaginary. \[x^4-81=0 \rightarrow x^4=81 \rightarrow x^2=\pm 9 \rightarrow x=\pm 3, \pm 3i\]